OpenStudy (anonymous):

Amy is twice as old as jenny. five years ago, amy was three times as old as jenny. find jenny's age.

OpenStudy (anonymous):

Let's first find variables for Amy and Jenny. A = Amy J = Jenny

OpenStudy (anonymous):

Ohhkaay

OpenStudy (anonymous):

A = 2J so A - 2J = 0 A - 5 = 3(J - 5) A - 5 = 3J - 15 + 5 A - 3J = - 10 Since A - 2J = 0 A - 3J = -10 So -J = -10; therefore, J = 10 Now we've found that Jenny is ten, and since Amy is twice as old: Jenny = 10, Amy = 20

OpenStudy (anonymous):

i get it now thanks